# Tradeoffs Between Alignment and Helpfulness in Language Models with Representation Engineering

### v4

@@ 3 Main Results
The proof is presented in appendix C and the assumption formally defined in appendix A. The above bound is illustrated in figure 3b for different values of $\beta$. As can be seen, around $r_{e}=0$, the bound is parabolic, i.e.  the decrease is proportional to $-r_{e}^{2}$, this can be obtained by expanding the bound near $r_{e}=0$. On the other hand, for large $r_{e}$, we see a decay to zero at a rate proportional to $r_{e}^{-2}$, this can be obtained by expanding the bound for large $r_{e}$. This can be extended to multi-token answers, by enforcing the above result on each decoding step of the generated answer, as explained in appendix I.
@@
Importantly, this demonstrates that for large $r_{e}$, the helpfulness decays to zero, hence representation engineering significantly harms the model’s overall performance, while for small $r_{e}$, it can initially decrease more slowly (parabolically) around $r_{e}=0$, hence the model’s performance is relatively unharmed. For the second statement to be feasible, the true helpfulness and the bound need to be close when no representation engineering is performed. The difference between the two at $r_{e}=0$ is bounded by $1-P_{0}$, such that for queries with high probability of being answered correctly without representation engineering, i.e. $P_{0}\approx 1$, the true helpfulness and the bound will be close, guaranteeing the parabolic bound to be meaningful.
@@
The parameter $\alpha\in[0,1]$ measures the tightness of the bound at $r_{e}=0$, since the true helpfulness at $r_{e}=0$ is $P_{0}$, while our helpfulness bound is $\frac{P_{0}}{P_{0}+\alpha(1-P_{0})}$. Thus $\alpha=1$ (and $\epsilon=0$) means the bound at $r_{e}=0$ coincides with the true helpfulness, while smaller $\alpha$ means the bound overshoots the true helpfulness. In our results, we obtain $\alpha\leq 0.5$. Figure 3 depicts this overshooting for $\alpha=0.25$. Even so, as explained above, the tightness is at least $1-P_{0}$ regardless of $\alpha$, so it is always meaningful for queries the model is initially helpful on.
@@
The product of parameters $\lambda\sigma\beta$ are a measure for the rate/curvature of the quadratic decay, as they are the coefficient multiplying $r_{e}^{2}$. $\lambda$ is the same scaling parameter from theorem 1, $\sigma$ represents the standard deviation of random noise added to the logits due to representation engineering (depicted in figure 3a and formally defined in A), and $\beta$ is the minimum between two weighted sums of positive variables with parameter $\sigma^{\prime}=1$. In section 4, we present our empirical estimation $\lambda\sigma\beta$ in the range $0.1-0.66$, both based on the logit noise condition and direct helpfulness measurement. Hence the decay should be felt at coefficients $r_{e}$ of order of size $1$.
@@
Lastly, when considering the average helpfulness over a dataset in a scenario where the number of answers is constant, $N$ (such as multiple choice questions), we obtain that on average, the model will converge to answering $1/N$ of the questions correctly as representation engineering is increased:
Corollary 2.
3.1 Interpretation of results – a tradeoff between alignment and usefulness
The combination of the two results show that alignment improves linearly with the norm of the vectors injected in representation engineering while helpfulness is decreased quadratically. This means that when injecting vectors of small norms, the improvement of alignment is initially faster than the decrease in helpfulness, possibly indicating that there is a regime where representation engineering is more effective. See figure 1 for an illustration of this intuition.
@@ 4 Empirical results
Here we will calculate alignment and helpfulness as defined above and observe how they change as we increase the vector norms of representation engineering. In principle, theorem 1 shows the dynamics of behavior flipping its signs due to representation engineering, thus to demonstrate it, we use representation engineering to show an increase in alignment of an unaligned pretrained model (specifically we use Llama 2 13B Touvron et al. (2023)), and a decrease in alignment of an aligned RLHF model (we use Llama 2 13B chat Touvron et al. (2023)). In appendix E, we perform the experiments on Llama 3.1 8B Dubey et al. (2024) as well. For the unaligned model, we calculate the behavior expectation w.r.t. behaviors “harmless" and “not-racist", as a function of representation engineering vector coefficients and show an improvement in alignment. For the aligned model, we do the same for the behaviors “harmful" and “racist" and show a decrease in alignment. Our experiments show an effect of representation engineering on alignment that matches theorem 1. Then, we calculate helpfulness as the probability of answering queries correctly when the model is injected with the same behavior altering vectors. Our experiments demonstrate that helpfulness changes as in theorem 2. Additional experimental details can be found in appendix E.
We follow the work of Zou et al. (2023a) to extract the vectors used in representation engineering: We use datasets comprised of pairs of positive and negative statements w.r.t. a behavior. The difference between the representations of the pairs are used to find latent space directions that can steer the model’s responses from negative to positive behaviors or vice versa. For the “harmful" behavior on the aligned model, we extracted harmful and unharmful instructions from AdvBench Robey et al. (2021; 2022) and shareGPT respectively. For “harmless" behavior on the the unaligned model, the above approach of contrasting positive and negative requests does not work, since the model did not undergo alignment, so it equally agrees to answer both types of requests. Thus, the produced engineered representations never steer the model towards not answering any request. So, inspired by the method of preference learning, we contrast aligned and misaligned responses to harmful instructions from AdvBench. For “racism" on the aligned model, we used biased statements from the StereoSet dataset Nadeem et al. (2020) followed by aligned and misaligned responses, and contrasted them. For “not-racist" on the unaligned model, we flipped the sign of the vectors to reverse the steering direction. The vectors obtained here were used to calculate behavior expectation and helpfulness of the model as the norm of the vectors increased.
@@
Figure 4a(c) shows harmless (harmful) behavior expectation as a function of harmless (harmful) PCA vector coefficients injected into the misaligned (aligned) model. Figure 4b(d) shows not-racist (racist) behavior expectation as a function of bias PCA vector coefficients injected into the misaligned (aligned) model. Overall we see that on both behaviors and both models, the behavior expectation changes like a hyperbolic tangent, as expected of theorem 1, which can be seen by the fitted curve of the data to a bound of the form of theorem 1 when using $\Delta\lambda$ as a free parameter that fits the measurements. The value of $\Delta\lambda$ corresponding to the curve is $0.5-3$ while our empirically estimated value of $\Delta\lambda$ from the data based on the linear classification condition of the last hidden layer change is $0.1-0.4$ (for details on the empirical estimation see appendix A.3). The difference between these two ranges may be attributed to the method of the empirical estimation of $\Delta$ that looks for an upper bound on it on the entire $r_{e}$ range, while the main change in alignment in figure 4 occurs in a more specific range, where the upper bound of $\Delta$ is evidently bigger. We note that for all behaviors, $r_{e}=2.5$ suffices for a significant change in behavior expectation, taking it from negative to positive. It is left to observe the decrease in helpfulness and verify that it is not too small.
@@
To calculate helpfulness, we tested the model on two tasks. The first is knowledge based question answering (MMLU), which allows a clean test for the single token theoretical results (theorem 2), and the second is coding tasks (HumanEval), which allows to verify the single token results persist for tasks with multiple-token answers. Importantly, we injected the model with the same vectors used to alter the model’s behavior in the alignment measurement. In appendix A.3 we show that while these vectors typically separate between aligned and misaligned responses in the model’s latent space when the prompt is related to the behavior in question, when prompted with knowledge based questions, the tokens enhanced are random with a trend that fits our theoretical assumptions.
First, we queried the model with multiple choice questions from the MMLU dataset Hendrycks et al. (2020) over a variety of domains (international law, medical genetics, high school computer science) and calculated the probability that the model assigns the correct one answer. This was measured as a function of injected vector coefficients inserted to the model for the behaviors above. Figure 5a(c) shows this for the harmless (harmful) behavior vectors on the misaligned (aligned) model and 5b(c) shows this for the not-racist (racist) behavior vectors for the misaligned (aligned) model. Here we restricted the probabilities to the answers of the multiple choice question, A,B,C,D. We also performed the experiment in a sampling setting, where we sampled full responses to the questions, and calculated the accuracy on the dataset, where similar trends were observed (see appendix E).
According to theorem 2, around $r_{e}=0$, helpfulness should decrease parabolically, which can be seen by Taylor expanding the bounds, yielding $P_{correct}\approx f(0)+\frac{1}{2}f^{\prime\prime}(0)r_{e}^{2}$ where $f^{\prime\prime}(0)<0$. Then, according to corollary 2, since there are $N=4$ answers to choose from, the probability of the correct answer should on average converge to $1/4$. To demonstrate this behavior in the empirical measurements, we plot a bound of the form of theorem 2 with the boundary conditions of corollary 2 (for further explanation on the theoretical justification of using this bound see appendix E.4). We do so with free parameter $\lambda\sigma\beta$, and $\alpha$ set to $1$ (in our theoretical bound it is smaller, but it is due to the bound being centered at $r_{e}=0$ while the peak is not guaranteed to be, hence it may overshoot), from which we find $\lambda\sigma\beta$ in the range of $0.33-0.66$. With our empirically estimated values of $\lambda\sigma\beta$ from direct measurement of the noise injected due to representation engineering, in the range $0.1-0.4$ (see details on the empirical estimation in appendix A.3).
@@
Notably, for $r_{e}=2.5$, the decrease in helpfulness is still not too great, while as mentioned previously, alignment is significantly increased. Further note that the decrease in helpfulness is not attributed to the model’s refusal to answer questions, as one might suspect for an injection of harmless vectors. This is because for both positive and negative coefficients the helpfulness drops, while the refusal to answer harmful queries grows only in one direction.
Next, we tested the model’s coding skills with the humaneval dataset. as can be seen in figure 6, the model’s performance is peaked around $r_{e}=0$, and it decays parabolically ar $r_{e}$ increases. We note that the asymmetry between positive and negative coefficients is captured in our theoretical bounds.
@@ 5 Discussion
In this work, we study the benefits of representation editing for LLM alignment from a theoretical perspective. We find that increasing the magnitude of the vectors injected to the model leads to improved alignment; we theoretically quantify this improvement as linear in the vectors’ magnitude, and validate our result empirically. A practical outcome of our result is a guarantee of alignment when using the representation engineering method. Such theoretical guarantees cannot be made without altering the model at inference time – Wolf et al. (2023) show that prompt based alignment methods can always be undone. Our result thus crystallizes an inherent advantage of representation engineering over competing alignment methods.
On the other hand, our framework indicates a degradation of the model’s general capabilities when representation engineering is applied. We theoretically quantify this degradation to be parabolic in the injected vectors’ magnitude, which puts a bound on the strength with which representation engineering should be performed to keep the model reliable for different uses. While our theoretical bound is an upper bound on the helpfulness, we observe this parabolic behavior empirically as well.
While representation engineering is an emerging field, editing interpretable features of models on the representation level in order to control them scales to SOTA models such as Sonnet’s Claude Templeton (2024). In principle, our framework may be generalized for theoretically analyzing the effects of normal finetuneing on alignment and helpfulness, as it too amounts to a change in the LLM representations to maximize the likelihood of desired outputs. In particular, each step in preference learning is equivalent to a representation injection with coefficient that equals to the learning rate (see appendix G). However, we leave this for future work, as finetuning creates small changes to the model’s representation at each training step on several behaviors, that sums to a large overall change, while representation engineering takes a large step in one behavioral direction. As a result, the change to the representations in a representation engineering process on one behavior creates random noise on the others (assumption 2), unlike a finetuning process where this does not necessarily happen. Hence in regards of maintaining helpfulness, finetuning has an advantage, however, representation engineering does enjoy the benefit of an online controllable step size in the desired behavior which allows to effectively manipulate the specific behavior at inference time.
Overall, we hope that our theoretical work will shed light on the mechanism of representation engineering, which constitutes a new interesting direction for language model alignment.
@@ A.1 Introduction of assumptions
Given a query $q$, the change to the last hidden layer of a model due to representation engineering, $\delta r_{e}(q)=r^{(L)}(q,r_{e})-r^{(L)}(q,0)$ , linearly classifies the representations of positive and negative answers to a query $q$ with margin $\Delta$, where the positive and negative answers are defined with respect to a behavior scoring function $B:\Sigma\rightarrow\{-1,+1\}$:
Note that the above assumption can be relaxed from a hard margin to a soft margin assumption, where $\delta r_{e}(q)$ classifies the representations of positive and negative answers, but part of the misaligned/aligned answers’ representations are misclassified as aligned/misaligned. This yields similar results to theorem 1 that are shown in appendix F.
For queries whose topic is unrelated to the behavior with respect to which representation engineering is performed, we expect the change to the last layer representation to be somewhat random on the highest probability tokens as they answer a question that is unrelated to the behavior whose vectors are injected to the model. Intuitively, the change to the final layer representation has no preference for a correct token over an incorrect token, so an incorrect answer is just as likely to be on one side or the other of the plane defined by the vertical $\delta r_{e}(q)$ that passes through the correct answer representation.
Assumption 2.
When sampling an answer to a query $q$ that is unrelated to the behavior of representation engineering, the vector $\delta r_{e}(q)=r^{(L)}(q,r_{e})-r^{(L)}(q,0)$, i.e., the resulting change to the last hidden layer representation due to the steering vectors from all layers, is random with the following coordinate-wise distribution on the $T$ highest probability tokens making $1-\epsilon$ of the probability mass:
Where $D$ is some continuous distribution with variance $\sigma^{2}>0$.
This defines a random directionality of $\delta r_{e}(q)$ w.r.t. the representations of answers. It is used in theorem 2 to formalize that representation engineering is a “perpendicular" direction to the query’s relevant answer representations.
Finally, we assume that for small coefficients of representation engineering $r_{e}$, the norm of the change to the last hidden layer representation is linear in $r_{e}$:
Assumption 3.
Let $P_{\theta,r_{e}}(\cdot|q)$ be a language model prompted with query $q$. The change to the last hidden layer representation due to representation engineering with coefficient $r_{e}$, denoted by $\delta r_{e}(q)=r^{(L)}(q,r_{e})-r^{(L)}(q,0)$ satisfies:

### v5

@@ 3 Main Results
The proof is presented in appendix C and the assumption formally defined in appendix A. The above bound is illustrated in figure 3b for different values of $\beta$. As can be seen, around $r_{e}=0$, the bound is parabolic, i.e.  the decrease is proportional to $-r_{e}^{2}$, obtained by expanding the bound near $r_{e}=0$. On the other hand, for large $r_{e}$, we see a decay to zero at a rate proportional to $r_{e}^{-2}$, obtained by expanding the bound for large $r_{e}$. This result can be extended to multi-token answers, by enforcing the above result on each decoding step of the generated answer, as explained in appendix K.
@@
Importantly, this demonstrates that while large vector injections harm the model’s overall performance, for small injections, the model’s performance is relatively unharmed due to the slow (parabolic) decrease with norm around $r_{e}=0$. For the second statement to be feasible, the true helpfulness and the bound need to be close when no steering is performed. Indeed, the difference between the two at $r_{e}=0$ is bounded by $1-P_{0}$, such that for queries with high probability of being answered correctly without steering, i.e. $P_{0}\approx 1$, the true helpfulness and the bound will be close, guaranteeing the parabolic bound to be meaningful.
@@
The parameter $\alpha\in[0,1]$ measures the tightness of the bound at $r_{e}=0$, since the true helpfulness at $r_{e}=0$ is $P_{0}$, while our helpfulness bound is $\frac{P_{0}}{P_{0}+\alpha(1-P_{0})}$. Thus $\alpha=1$ (and $\epsilon=0$) means the bound at $r_{e}=0$ coincides with the true helpfulness, while smaller $\alpha$ means the bound overshoots it. In our results, we obtain $\alpha\leq 0.5$. Figure 3 depicts this overshooting for $\alpha=0.25$. Even so, as explained above, the tightness is at least $1-P_{0}$ regardless of $\alpha$, so it is always meaningful for queries the model is initially helpful on.
@@
The product of parameters $\lambda\sigma\beta$ measures the rate/curvature of the quadratic decay, as they are the coefficient multiplying $r_{e}^{2}$. $\lambda$ is the same scaling parameter from theorem 1, $\sigma$ is the standard deviation of random noise added to the logits due to representation engineering (depicted in figure 3a and formally defined in A). $\beta$ is the minimum between two weighted sums of positive variables with parameter $\sigma^{\prime}=1$. In section 4, we present an empirical estimation for $\lambda\sigma\beta$ in the range $0.1-0.66$, based on the logit noise condition and direct helpfulness measurement. Hence the decay becomes strong at coefficients $r_{e}$ of size $1-10$.
@@
A tradeoff between alignment and usefulness:
The combination of the two results shows alignment improves linearly with the norm of the steering vectors while helpfulness is decreased quadratically. This means that when injecting vectors of small norms, the improvement of alignment is initially faster than the decrease in helpfulness, indicating a regime where steering is more effective. See figure 1 for an illustration of this intuition.
@@ 4 Empirical results
Here we will calculate alignment and helpfulness as defined above and observe how they are affected by increasing norms of steering vectors. Theorem 1 shows how alignment can increase/decrease due to steering, thus to demonstrate it, we increase the alignment of an unaligned pretrained model w.r.t.  “harmless" and “not-racist" behaviors (specifically we use Llama 2 13B (Touvron et al., 2023)), and conversely, misalign an aligned RLHF model w.r.t. “harmful" and “racist" behaviors (Llama 2 13B chat (Touvron et al., 2023)). Then, we calculate helpfulness as the probability of answering queries correctly when the model is injected with the same behavior altering vectors. The experiments show an effect on alignment matching theorem 1 and on helpfulenss matching theorem 2. Additional experimental details can be found in appendix F as well as results for Llama 3.1 8B (Dubey et al., 2024). We note the goal of the experiments is to demonstrate the theoretical bounds showing an enhancement of alignment with a concept and a helpfulness decrease due to steering, and that a complementary experimental demonstration of these with more behaviors is shown on Claude 3 Sonnet with social biases when using feature steering (Durmus et al., ).
We follow the work of Zou et al. (2023a) to extract the vectors used in representation engineering: Pairs of positive and negative statements w.r.t. a behavior, are forward passed through the model, and the differences between representations of the pairs are used to find latent space directions that steer the model’s responses from negative to positive behaviors or vice versa. For the “harmful" behavior on the aligned model, we extracted harmful and unharmful instructions from AdvBench (Robey et al., 2021; 2022) and shareGPT respectively. For “harmless" behavior on the unaligned model, the approach of contrasting positive and negative requests does not work, as the model agrees to answer both types of requests, so contrasting them does not steer the model towards not answering a request. Instead, inspired by the method of preference learning, we contrast aligned and misaligned responses to harmful instructions from AdvBench. For “racism" on the aligned model, we used biased and unbiased statements from the StereoSet dataset (Nadeem et al., 2020). For “not-racist" on the unaligned model, we used the racist statements from above, followed by aligned and misaligned responses. The obtained vectors were used to calculate behavior expectation and helpfulness of the model as the norm of the vectors increased.
@@
Figure 4 shows behavior expectation as a function of corresponding PCA vector coefficients injected into the models. Overall we see that on both behaviors and both models, the behavior expectation changes like a hyperbolic tangent, as expected of theorem 1, which can be seen by the fitted curve of the data to a bound of the form of theorem 1 when using $\Delta\lambda$ as a free parameter that fits the measurements. The value of $\Delta\lambda$ corresponding to the curve is $0.5-3$ while our empirically estimated value of $\Delta\lambda$ from the data based on the linear classification condition of the last hidden layer change is $0.1-0.4$ (for details and explanation for these differences see appendix A.3). We note that for all behaviors, $r_{e}=2.5$ suffices for a significant change in behavior expectation, taking it from negative to positive. It is left to observe the decrease in helpfulness and verify that it is not too big.
@@
To calculate helpfulness, we tested the model on two tasks. The first is knowledge based question answering, for a clean test of the single token theoretical results (theorem 2). The second is code generation, to verify the single token results persist for tasks with multiple-token answers. Importantly, we injected the model with the same vectors used to alter the model’s behavior in the alignment measurement.
For the first task, we queried the model with multiple choice questions from the MMLU dataset (Hendrycks et al., 2020) over a variety of domains (e.g. international law, medical genetics) and calculated the probability that the model assigns the correct answer. This was done both by calculating the probabilities of the multiple choice answers, A,B,C,D, and in appendix F by sampling full responses to the questions and measuring the accuracy, yielding similar results. This was measured as a function of injected vector coefficients inserted to the model for the behaviors above. Figure 5 shows the results for the different behaviors and models. We plot a bound of the form of theorem 2 to demonstrate the predicted parabolic behavior. We do so with free parameter $\lambda\sigma\beta$ from which we find $\lambda\sigma\beta$ in the range of $0.33$ to $0.66$ (see appendix F.4). This is in accordance with our empirically estimated values of $0.1$ to $0.4$ for $\lambda\sigma\beta$ from direct measurement of the noise injected to the model due to representation engineering in appendix A.3. Notably, for $r_{e}=2.5$, the decrease in helpfulness is still not too great, while as mentioned previously, alignment is significantly increased.
@@
For the second task, we tested the model’s coding skills with the humaneval dataset (Chen et al., 2021). We present the results in appendix G. The model’s performance is peaked around $r_{e}=0$, and it decays parabolically ar $r_{e}$ increases, as predicted in theorem 2.
@@ 5 Discussion
In this work, we study the benefits of steering methods for LLM alignment from a theoretical perspective. We find that increasing the magnitude of the vectors injected to the model leads to improved alignment; we theoretically quantify this improvement as linear in the vectors’ magnitude, and validate our result empirically. A practical outcome of our result is a guarantee of alignment when using the representation engineering method. Such theoretical guarantees cannot be made without altering the model at inference time – Wolf et al. (2023) show that prompt based alignment methods can always be undone. Our result thus crystallizes an inherent advantage of steering over competing alignment methods.
On the other hand, our framework indicates a degradation of the model’s general capabilities when steering is applied. We theoretically quantify this degradation to be parabolic in the injected vectors’ magnitude, which puts a bound on the strength with which steering should be performed to keep the model reliable for different uses. While our theoretical bound is an upper bound on the helpfulness, we observe this parabolic behavior empirically as well.
While steering is an emerging field, editing interpretable features of models on the representation level in order to control them scales to SOTA models such as Anthropic’s Claude 3 Sonnet (Templeton, 2024; Durmus et al., ). In principle, our framework may be generalized for theoretically analyzing the effects of normal finetuneing on alignment and helpfulness, as it too amounts to a change in the LLM representations to maximize the likelihood of desired outputs. In particular, each step in preference learning is equivalent to steering with coefficient that equals to the learning rate (see appendix I), and indeed similar tradeoffs have been observed for finetuning (Tan et al., 2024). However, we leave this for future work, as finetuning creates small changes to the model’s representation at each training step on several behaviors, that sums to a large overall change, while steering takes a large step in one direction. As a result, the change to the representations in a steering process on one behavior creates random noise on the others (assumption 3), unlike a finetuning process where this does not necessarily happen. Hence in regards of maintaining helpfulness, finetuning has an advantage, however, steering does enjoy the benefit of an online controllable step size in the desired behavior for effective manipulation at inference time.
Overall, we hope that our theoretical work will shed light on the mechanism of steering, which constitutes a new interesting direction for language model alignment.
@@ A.1 Introduction of assumptions
We assume that for small coefficients of representation steering $r_{e}$, the norm of the change to the last hidden layer representation is linear in $r_{e}$:
@@
Let $P_{\theta,r_{e}}(\cdot|q)$ be a language model prompted with query $q$. The change to the last hidden layer representation due to steering with coefficient $r_{e}$, denoted by $\delta r_{e}(q)=r^{(L)}(q,r_{e})-r^{(L)}(q,0)$ satisfies:
@@
Assumption 2.
Given a query $q$, the change to the last hidden layer of a model due to steering, $\delta r_{e}(q)=r^{(L)}(q,r_{e})-r^{(L)}(q,0)$ , linearly classifies the representations of positive and negative answers to a query $q$ with margin $\Delta$, where the positive and negative answers are defined with respect to a behavior scoring function $B:\Sigma\rightarrow\{-1,+1\}$:
Note that the above assumption can be relaxed from a hard margin to a soft margin assumption, where $\delta r_{e}(q)$ classifies the representations of positive and negative answers, but part of the misaligned/aligned answers’ representations are misclassified as aligned/misaligned. This yields similar results to theorem 1 that are shown in appendix H.
For queries whose topic is unrelated to the behavior with respect to which steering is performed, we expect the change to the last layer representation to be somewhat random on the highest probability tokens as they answer a question that is unrelated to the behavior whose vectors are injected to the model. Intuitively, the change to the final layer representation has no preference for a correct token over an incorrect token, so an incorrect answer is just as likely to be on one side or the other of the plane defined by the vertical $\delta r_{e}(q)$ that passes through the correct answer representation.
Assumption 3.
When sampling an answer to a query $q$ that is unrelated to the behavior of steering, the vector $\delta r_{e}(q)=r^{(L)}(q,r_{e})-r^{(L)}(q,0)$, i.e., the resulting change to the last hidden layer representation due to the steering vectors from all layers, is random with the following coordinate-wise distribution on the $T$ highest probability tokens making $1-\epsilon$ of the probability mass:
Where $D$ is some continuous distribution with variance $\sigma^{2}>0$.
This defines a random directionality of $\delta r_{e}(q)$ w.r.t. the representations of answers. It is used in theorem 2 to formalize that steering is a “perpendicular" direction to the query’s relevant answer representations.